OpenStudy (anonymous):

Can somebody please explain how you get the very last step of this equation? Length of a line segment.

OpenStudy (anonymous):

Here it is.

OpenStudy (anonymous):

\[152=9\times 17\]

OpenStudy (anonymous):

oops i meant \(153=9\times 17\)

OpenStudy (anonymous):

and so \(\sqrt{153}=\sqrt{9\times 17}=\sqrt9\sqrt{17}=3\sqrt{17}\)

OpenStudy (anonymous):

Ok I see, thanks!